OpenStudy (anonymous):

Okay. I am so dumb at math. Goodness. I get A's in everything else, but I am so not getting the computation steps to get an answer. So here is the problem. What is 7 percent sales tax on $279? I know that I have to multiply 0.7 X 279, but is the answer $19.53 tax or is it $1.95? Please help!

OpenStudy (anonymous):

you multiply 0.07 not 0.7, 0.7 would be 70% sales tax

OpenStudy (anonymous):

try using fractions 7% = 7/100 times 279 or 70% = 70/100 times 279

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

BONK

OpenStudy (anonymous):

how then you got